OpenStudy (anonymous):

What Is 1. x+10=-12, 2. y-12=15, 3. 2x=-18, 4. x/3= -4, 5. 3y=15

OpenStudy (anonymous):

okay I will start with number one to get 10 on the other side you have to subtract\[x + 10 - 10 = -12 - 10\]

OpenStudy (anonymous):

Is -2 The Answer?

OpenStudy (anonymous):

what is -12 - 10

OpenStudy (anonymous):

-22?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

2. y-12=15 now to do this one you want to add 12 to both sides

OpenStudy (anonymous):

y=37 Right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

3. 2x=-18 this one you want to divide what do you think you are going to divide

OpenStudy (anonymous):

2 On Each Side? Right?

OpenStudy (anonymous):

And The Answer Would Come Out To Be -9?

OpenStudy (anonymous):

right now do you think you see what to do with the rest

OpenStudy (anonymous):

Yes! But I Dont Get Number 4

OpenStudy (anonymous):

You Multiply 3*4 Right And Come Up With x=-12 Right?

OpenStudy (anonymous):

right just watch your signs

OpenStudy (anonymous):

I Thought If You Mulitiply 3*-4 It Will Be A Negative Cause' Of The Rules!

OpenStudy (anonymous):

right but the why you wrote it the first time you forgot to but the - sign but the answer is -12

OpenStudy (anonymous):

Oh! Sorry Thanks I Asked All My Questions Now!